Create a bucket by rotating around the y axis the curve y=2ln(x−4)y=2ln(x-4) from y = 0 to y = 4. If this bucket contains a liquid with density 820 kg/m3 to a height of 3 meters, find the work required to pump liquid out of this bucket (over the top edge). Use 9.8 m/s2 for gravity
Answer #1

को 1-4 J= 22n (x-4) 7x X-5 P = 820 kyms om = pdV dM :. .. dW = PA(Y) 49 (M)7 (५-५) (५-५) PA(५) dy twa 3 : W = हो (५-५) A(s) lPage No. Aly) = Tx? 3 W = pg (4-4) T x² dy yo $14_Y) y = 2 ln (x-4 dy = z de sty ute W =28 g (4=2 lmfx=W) x2 dx x-Y W = 20gf